OpenStudy (anonymous):

Find the 15th term of the arithmetic sequence. 2, 3.5, 5, 6.5, ...

OpenStudy (anonymous):

If we needed the 12th term, we would calculate a12 = (0.01)612 - 1 = (0.01)611 ... Given our generic arithmetic sequence a1, a2, a3, a4, ..., an, ... add the first 10 terms of sequence B, we would know that a1 = 0.01, r = 6 and n = 10.

OpenStudy (anonymous):

The difference in each term is d=1.5. To find the nth term: a(n)=a(1)+(n-1)d

OpenStudy (anonymous):

a(15)=2+14(1.5)

OpenStudy (anonymous):

so it would be 21.5?

OpenStudy (anonymous):

no...evaluate the equation i just wrote